# Price formation without fuel costs: the interaction of elastic demand with storage bidding

### v1

@@ 4.5 Cost recovery
Figure 7: Cost recovery of each component as a fraction of investment, depending on demand modeling approach, operational foresight and capacity perturbations for 35-year LT plus 35-year ST optimizations.
Cost recovery refers to the fraction of the investment and operation costs that are recovered by the revenues from the operation of the assets. Fig. 7 shows the cost recovery for each asset in different scenarios. For the LT model, the cost recovery is perfect for all components, following the zero-profit rule shown in Brown and Reichenberg [47].
Tranitioning from the LT to the ST model with perfect foresight, drastic differences are observed between demand modelling cases. The high overshoot of cost recovery rates for the inelastic case, due to the high prices seen in Fig. 6, is not observed in the elastic case and highlights the inadequacy of modelling with perfectly inelastic demand.
For scenarios with myopic operational foresight in unseen years, the differences between inelastic and elastic demand cases are pruned by the heuristic storage bids. Especially for the elastic case, all components exceed or come close to full cost recovery. The lowest cost recovery is 97.2% for the electrolyzer. With over-capacities of 5%, cost recovery falls short and varies between 67% and 82%, incentivizing market exits. The opposite holds for deficient capacities, where profits between 14% and 190% incentivize certain market entries.
All results shown in preceding Sections 4.1, 4.2, 4.3, 4.4 and 4.5 for Germany are consistent with those observed for a more solar-dominated system in Spain and a more wind-dominated system in the United Kingdom, as shown in D.
@@ 5.1 Consquences for interpreting prices in energy system models
Our results establish that the strong price bifurcations seen in many capacity expansion models are due to modeling demand without price elasticity. The combination of step-like inelastic demand curves with step-like supply curves from VRE together cause the price bifurcations. Given that today’s electricity markets already feature price-sensitive demand [43], price duration curves from these models are unrealistic and should not be taken at face value.
While average prices and market values might be indicative in models with inelastic demand, the shape of the price duration curves is not stable as the number of optimized periods is changed, nor under transition from long-term to short-term modeling, nor under small perturbations of capacity, nor between different weather years. Prices either collapse to zero for extended periods or spike dramatically. The concerns in the literature, for instance, that revenues of generation and storage would be realized in just very few hours [2, 3, 4, 5], are rooted in these observations but are, in fact, interpretations of a modeling artifact.
The short-term elasticity level observed in the German wholesale electricity market today (cf. [43]) would already be sufficient to stabilize price formation and yield a completely different picture about when revenues are made. The price formation with elastic demand is then identical between long- and short-term models, stable under capacity perturbations, stable between between weather years, and allows for more consistent and evenly spread cost recovery.
The prospect of more cross-sectoral integration further opens new opportunities for demand-side management of electric vehicles, heat pumps, and industrial processes, beyond today’s elasticity level. We find that, especially if there is no outside mechanism to stabilize fuel prices, such demand elasticity is crucial for non-binary price formation.
5.2 Relating the modeling results to real electricity markets
@@
Cost recovery is naturally imperfect with imperfect information. For instance, with capacities below or above the long-term equilibrium, there would be incentives for market entries or exits as asset owners would adjust their capacities to prevailing market conditions. The more regular and lower high price episodes in the elastic case would provide investors with a much more stable investment environment than the infrequent extreme high price episodes seen with the inelastic case, thus helping to avoid boom and bust cycles. The political tension between allowing scarcity pricing and remunerating dispatchable generators via capacity markets to guarantee security of supply is present in our high-VRE scenarios, just as it is in today’s markets.
Our results suggest that the energy-only market remains functional in future energy systems with high VRE penetration.
Although our results suggest that the energy-only market would remain functional in future energy systems with high VRE penetration, it may not be sufficient on its own to attain a low-cost system. To lower investor risk and reduce financing costs, support guarantees such as investment subsidies, contracts for difference or feed-in tariffs may still be advisable [55]. The strike prices of these contracts could use the energy-only market for orientation towards efficient prices.
The assumption of a constant marginal storage value for the myopic dispatch is strong, and in reality, storage operators would adjust their bidding based on their filling state and medium to long-range weather forecasts [35]. Moreover, a flat price would mean storage operators could not perform arbitrage and recover their costs. Potentially, long-term contracts for industrial offtake or global supply (e.g. H2Global [56]) could stabilize the hydrogen price.
@@
5.3 Comparison of chemical storage with other current fuels
In terms of their impact on price formation in electricity markets, we can compare chemical storage to other currently used fuels. For instance, hydroelectricity shares many similarities with chemical storage. Both can provide long-duration storage that stabilizes prices through opportunity. However, unlike chemical storage, hydroelectric storage is constrained by natural inflow, self-discharge, and the fact that the storage medium cannot be traded on global markets. Chemical storage expansion is not geologically limited and offers more control over dimensioning and scaling.
The price of fossil fuels like gas and coal also stems from a mix of production and opportunity costs, which is similar to how the value of chemical storage is determined in our model. For hydrogen, production costs include wind and solar generation and power-to-X elements. For fossil fuels, it is the energy and labor cost for extraction, which may be further impacted by emission prices or scarcity prices due to supply disruptions. Similarly, prices are linked to feedstock availability for electricity production from biomass and the mining, enrichment, and fabrication of uranium into fuel rods for nuclear power. The connection to currently used fuels becomes even more apparent when considering green fuel imports for power backup, where the price is decoupled from the domestic electricity system. The price formation mechanism is identical to that of a fossil-based system.
@@ 6 Conclusion
Our study has addressed several contradictions seen in the literature about the viability of electricity markets at high penetrations of wind and solar. Our results demonstrate that the price structure of many zero- and high-price hours is a symptom of capacity expansion models using inelastic demand. Already a slightly elastic demand alleviates these symptoms and reduces the fraction of zero-price hours from 90% to around 30%. Besides eliminating price bifurcations, modelling demand elasticity stabilizes prices between weather years, which reduces revenue risks, and aligns prices observed in long- and short-term models. Therefore, we conclude that electricity models using inelastic demand are inadequate for analyzing price formation at high penetrations of variable renewable energy.
Moreover, elastic demand not only stabilizes price formation but also facilitates the operation of long-term storage with myopic foresight. The stability of the marginal values of chemical storage, which are closely linked to electricity price formation, enables the development of storage bidding strategies. Even with the simple assumption of constant storage bids, we can achieve adequate system dispatch with similar prices, reliability metrics, and cost recovery levels, with only a few days of foresight, and even if capacities do not exactly match the long-term equilibrium.
With these insights, we address many of the concerns about price formation in energy-only markets with high penetrations of wind and solar. We conclude that the energy-only market can still play a key role in coordinating future dispatch and investment with current price elasticity levels.
@@
The open source code to reproduce our experiments, including the download and installation of data and software dependencies, is available on GitHub licensed under the MIT license at https://github.com/fneum/price-formation (v0.1.0). The results data was deposited on Zenodo under the CC-BY-4.0 license at https://doi.org/10.5281/zenodo.12759248 (v0.1.0).
@@ Acknowledgements
We thank Lina Reichenberg, Wolf-Peter Schill, Falko Ueckerdt, Robert Pietzcker, Gunnar Luderer, Chris Gong, Oliver Ruhnau, Richard Schmalensee, and Dharik Mallapragada for fruitful discussions.
@@ Appendix A Substitutions to improve numerics
Many of our initial runs with VOLL or elastic demand represented as outlined in Eq. 1 were numerically unstable and did not converge when many weather years were included. We found that making the substitutions $d_{c}\rightarrow D_{c}-g_{c}$ for VOLL and $d_{c}\rightarrow a_{c}/b_{c}-g_{c}$ for elastic demand drastically improved the convergence of the model while yielding the identical results, as they remove large constants from the objective.
The suggested substitutions switch from modeling the demand as a decision variable to modeling load shedding from fixed peak consumption ($D_{c}$ or $a_{c}/b_{c}$) as a decision variable. In terms of the objective function, it yields for VOLL
@@
Since the constant terms $V_{c}D_{c}$ and $a_{c}^{2}/2b_{c}$, which represent the welfare under the demand curve, do not affect the optimization, we can drop them from the objective function. What remains is a load-shedding generator with the cost curve $C_{c}(g_{c})=V_{c}g_{c}$ or $C_{c}(g_{c})=\frac{b_{c}}{2}g_{c}^{2}$ for the elastic case. In addition, a fixed demand of $D_{c}$ or $a_{c}/b_{c}$, respectively, needs to be added to the model.
This substitution trick can also be generalized for piecewise-linear demand curves with segments $c$, where the demand curves are not zero at $a_{c}/b_{c}$. We now apply the more general substitution $d_{c}\rightarrow D_{c}-g_{c}$ for each segment $c$.
@@
Appendix B Technology data
@@
H2 fuel cell / turbine 1164.0 €/kWel 5.00 %/a 50.0% 10 a
@@
Appendix C Additional figures

### v2

@@ 4.6 Cost recovery
Figure 7: Cost recovery of each component as a fraction of investment, depending on demand modelling approach, operational foresight and capacity perturbations for 35-year LT plus 35-year ST optimisations.
Cost recovery refers to the fraction of the investment and operation costs that are recovered by the revenues from the operation of the assets. Fig. 7 shows the cost recovery for each asset in different scenarios. For the LT model, the cost recovery is perfect for all components, following the zero-profit rule shown in Brown and Reichenberg [51]. The variability of revenue from year to year is, however, much higher for the VOLL case than the PWL-elastic case (Fig. S10).
Transitioning from the LT to the ST model with perfect foresight, drastic differences are observed between demand modelling cases. The high overshoot of cost recovery rates for the perfectly inelastic case until VOLL, due to the high prices seen in Fig. 6, is not observed in the elastic case and highlights the inadequacy of modelling with perfectly inelastic demand.
For scenarios with myopic operational foresight in unseen years, the differences between inelastic and elastic demand cases are pruned by the heuristic storage bids. Especially for the elastic case, all components exceed or come close to full cost recovery. The lowest cost recovery is 97.2% for the electrolyzer. With over-capacities of 5%, cost recovery falls short and varies between 67% and 82%, incentivising market exits. The opposite holds for deficient capacities, where profits between 14% and 190% incentivise certain market entries.
All results shown in preceding Sections 4.1, 4.2, 4.3, 4.4, 4.5 and 4.6 for Germany are consistent with those observed for a more solar-dominated system in Spain and a more wind-dominated system in the United Kingdom, as shown in F.
@@ 5.1 Consquences for interpreting prices in energy system models
Our results establish that the strong price bifurcations seen in many capacity expansion models are due to modelling demand without price elasticity. The combination of step-like inelastic demand curves with step-like supply curves from VRE together cause the price bifurcations. Given that today’s electricity markets already feature price-sensitive demand [47], price duration curves from these models are unrealistic and should not be taken at face value.
While average prices and market values might be indicative in models with perfectly inelastic demand, the shape of the price duration curves is not stable as the number of optimised periods is changed, nor under transition from long-term to short-term modelling, nor under small perturbations of capacity, nor between different weather years. Prices either collapse to zero for extended periods or spike dramatically. The concerns in the literature, for instance, that revenues of generation and storage would be realised in just very few hours [3, 4, 5, 6], are rooted in these observations but are, in fact, interpretations of a modelling artifact.
The short-term elasticity level observed in the German wholesale electricity market today (cf. [47]) would already be sufficient to stabilise price formation and yield a completely different picture about when revenues are made. The price formation with demand elasticity is then identical between long- and short-term models, stable under capacity perturbations, stable between between weather years, and allows for more consistent and evenly spread cost recovery.
The reason demand elasticity helps to stabilise prices is that it gives the storage a wide range of willingness-to-pay (WTP) values against which to calibrate its opportunity costs and arbitrage. With perfectly inelastic demand, the WTP jumps from zero to the value of lost load with nothing in between for the storage to bid against.
The prospect of more cross-sectoral integration further opens new opportunities for demand-side management (DSM) of electric vehicles, heat pumps, and industrial processes, beyond today’s elasticity level, although inter-temporal dependencies need to be treated with care. Modelling DSM with virtual storage units is not sufficient to resolve pricing problems, as the examples above with storage show; sloped demand curves with smoothly varying WTP are what is needed for stable price formation.
5.2 Relating the modelling results to real electricity markets
@@
Cost recovery is naturally imperfect with imperfect information. For instance, with capacities below or above the long-term equilibrium, there would be incentives for market entries or exits as asset owners would adjust their capacities to prevailing market conditions. The more regular and lower high price episodes in the elastic case would provide investors with a much more stable investment environment than the infrequent extreme high price episodes seen with the perfectly inelastic case, thus helping to avoid boom and bust cycles. The political tension between allowing scarcity pricing and remunerating dispatchable generators via capacity markets to guarantee security of supply is present in our high-VRE scenarios, just as it is in today’s markets.
Our results suggest that the energy-only market remains functional in future energy systems with high VRE penetration. However, it may not be sufficient on its own to attain a low-cost renewable system. Investor risk remains thanks to uncertainties surrounding inter-annual revenue variability, changing weather patterns due to climate change, other factors that can confound price forecasts (price cannibalisation from other renewables, CO2 price uncertainty) as well as regulatory risks and political risks from changing climate targets. To lower investor risk and reduce financing costs, support guarantees such as investment subsidies, contracts for difference or feed-in tariffs may still be advisable [63]. The strike prices of these contracts could use the energy-only market for orientation towards efficient prices.
The assumption of a constant marginal storage value for the myopic dispatch is strong, and in reality, storage operators would adjust their bidding based on their filling state and medium to long-range weather forecasts [38]. Moreover, a flat price would mean storage operators could not perform arbitrage and recover their costs. Potentially, long-term contracts for industrial offtake or global supply (e.g. H2Global [64]) could stabilise the hydrogen price.
@@
5.3 Comparison of chemical storage to conventional generation
In terms of their impact on price formation in electricity markets, we can compare chemical storage to conventional generation technologies. For instance, hydroelectricity shares many similarities with chemical storage. Both can provide long-duration storage that stabilises prices through long-term opportunity costs. However, unlike chemical storage, hydroelectric storage is constrained by natural inflow, self-discharge, and the fact that the storage medium cannot be traded on global markets. Hydroelectric inflow is easier to forecast on a seasonal basis that wind power fluctuations, thanks to predictable snow and rainfall patterns. On the other hand, chemical storage expansion is not geologically limited and offers more control over dimensioning and scaling.
The price of fossil fuels like gas and coal also stems from a mix of production and opportunity costs, which is similar to how the value of chemical storage is determined in our model. For hydrogen, production costs include wind and solar generation and power-to-X elements. For fossil fuels, it is land rents as well as the energy and labor costs for extraction, which may be further impacted by emission prices or scarcity prices due to supply disruptions. Similarly, prices are linked to feedstock availability for electricity production from biomass and the mining, enrichment, and fabrication of uranium into fuel rods for nuclear power. The connection to currently used fuels becomes even more apparent when considering green fuel imports for power backup, where the price is decoupled from the domestic electricity system. The price formation mechanism is then similar to that of a fossil-based system.
@@ 6 Conclusion
Our study addresses several contradictions seen in the literature about the viability of electricity markets at high penetrations of wind and solar. Our results demonstrate that price structures consisting of many zero- and high-price hours are a symptom of capacity expansion models using perfectly inelastic demand. Already a slightly elastic demand alleviates these symptoms and reduces the fraction of zero-price hours from 90% to around 30%. Besides eliminating price bifurcations, modelling demand elasticity stabilises prices between weather years, which reduces revenue risks, and aligns prices observed in long- and short-term models. Therefore, we conclude that electricity models using perfectly inelastic demand are inadequate for analysing price formation at high penetrations of variable renewable energy.
Demand elasticity not only stabilises price formation but also facilitates the operation of long-term storage with myopic foresight. The stability of the marginal values of chemical storage, which are closely linked to electricity price formation, enables the development of storage bidding strategies. Even with the simple assumption of constant storage bids, we can achieve adequate system dispatch with similar prices, reliability metrics, and cost recovery levels, with only a few days of foresight, and even if capacities do not exactly match the long-term equilibrium.
With these insights, we address many of the concerns about price formation in energy-only markets with high penetrations of wind and solar. We conclude that the energy-only market can still play a key role in coordinating future dispatch and investment given current price elasticity levels. If demand can be further flexibilised in future, this will only help to stabilise prices even more.
@@
The open source code to reproduce our experiments, including the download and installation of data and software dependencies, is available on GitHub licensed under the MIT license at https://github.com/fneum/price-formation (v0.2.0). The results data was deposited on Zenodo under the CC-BY-4.0 license at https://doi.org/10.5281/zenodo.12759247 (v0.2.0).
@@ Acknowledgements
We thank Lina Reichenberg, Wolf-Peter Schill, Falko Ueckerdt, Robert Pietzcker, Gunnar Luderer, Chris Gong, Oliver Ruhnau, Richard Schmalensee, Dharik Mallapragada, Emil Dimanchev, Richard Green and Tim Schittekatte for fruitful discussions.
@@ Appendix A Substitutions to improve numerics
Many of our initial runs with VOLL or demand elasticity represented as outlined in Eq. 1 were numerically unstable and did not converge when many weather years were included. We found that making the substitutions $d_{c}\rightarrow D_{c}-g_{c}$ for VOLL and $d_{c}\rightarrow a_{c}/b_{c}-g_{c}$ for demand elasticity drastically improved the convergence of the model while yielding the identical results, as they remove large constants from the objective.
The suggested substitutions switch from modelling the demand as a decision variable to modelling load shedding from fixed peak consumption ($D_{c}$ or $a_{c}/b_{c}$) as a decision variable. In terms of the objective function, it yields for VOLL
@@
Since the constant terms $V_{c}D_{c}$ and $a_{c}^{2}/2b_{c}$, which represent the welfare under the demand curve, do not affect the optimisation, we can drop them from the objective function. What remains is a load-shedding generator with the cost curve $C_{c}(g_{c})=V_{c}g_{c}$ or $C_{c}(g_{c})=\frac{b_{c}}{2}g_{c}^{2}$ for the elastic case. In addition, a fixed demand of $D_{c}$ or $a_{c}/b_{c}$, respectively, needs to be added to the model.
This substitution trick can also be generalised for piecewise-linear demand curves with segments $c$, where the demand curves are not zero at $a_{c}/b_{c}$. We now apply the more general substitution $d_{c}\rightarrow D_{c}-g_{c}$ for each segment $c$.
@@
Appendix B Including cross-elasticity terms
To account for how demand changes in one hour affect the demand curve in other hours, we extend the quadratic utility term in the objective function to include cross-elasticity terms:
$U_{c}(d_{c,t})=\sum_{t}\left(a_{c}d_{c,t}-\frac{b_{c}}{2}d_{c,t}^{2}+\sum_{k}\frac{\gamma_{c,t,k}}{2}d_{c,t}d_{c,k}\right)$ (15)
where $\gamma_{c,t,k}$ represents the cross-elasticity between hours $t$ and $k$ for segment $c$. The factor $\frac{1}{2}$ appears because each cross-term appears twice in the summation. The cross-elasticity coefficient is symmetric ($\gamma_{c,t,k}=\gamma_{c,k,t}$) and is only non-zero for adjacent hours within a specified time window:
$\gamma_{c,t,k}=\begin{cases}\gamma_{c}&\text{if }|t-k|\leq\hat{T}\text{ and }t\neq k\\ 0&\text{otherwise}\end{cases}$ (16)
where $\hat{T}$ is the maximum difference in time steps for considering cross-elasticity effects and $\gamma_{c}$ is chosen proportionally to $b_{c}$. For each demand segment $c$, we use $\gamma_{c}=\frac{b_{c}}{16}$ in the default case to ensure cross-effects do not dominate own-price elasticity. We also choose as default $\hat{T}=4$ hours. The diagonal elements ($t=k$) are zero since own-price effects are already captured by the quadratic term with coefficient $b_{c}$.
This approach means that if demand in hour $t$ decreases by 1 MW, the willingness to pay in all hours $k$ where $|t-k|\leq\hat{T}$ increases by $\gamma_{c}$/MWh, reflecting the increased need to consume in those hours to maintain the same total consumption level. The cross-elasticity does not change the price elasticity itself; the demand curve shifts up in those hours, but its slope (elasticity) stays the same. The objective function remains concave with the inclusion of cross-elasticity terms as long as $\gamma_{c}$ is small enough, such that we do not expect a strong computational impact.
For numerical stability, we apply the same substitution as in A, replacing $d_{c,t}\rightarrow D_{c}-g_{c,t}$ where $D_{c}=a_{c}/b_{c}$. For the cross-elastic terms, this yields:
$\sum_{k}\left(\frac{\gamma_{c,t,k}}{2}d_{c,t}d_{c,k}\right)$ $=\sum_{k}\left(\frac{\gamma_{c,t,k}}{2}(D_{c}-g_{c,t})(D_{c}-g_{c,k})\right)$
$=\sum_{k}\left(\frac{\gamma_{c,t,k}}{2}(D_{c}^{2}-D_{c}g_{c,t}-D_{c}g_{c,k}+g_{c,t}g_{c,k})\right)$
$=\sum_{k}\left(\frac{D_{c}^{2}\gamma_{c,t,k}}{2}-\frac{D_{c}\gamma_{c,t,k}}{2}g_{c,t}-\frac{D_{c}\gamma_{c,t,k}}{2}g_{c,k}+\frac{\gamma_{c,t,k}}{2}g_{c,t}g_{c,k}\right)$ (17)
The constant terms $\sum_{k}D_{c}^{2}\gamma_{c,t,k}/2$ can be dropped from the objective function. The remaining terms are added to the objective function as cost terms of load shedding generators $g_{c,t}$ for cost minimisation, analogous to A:
$\sum_{k}\left(\frac{D_{c}\gamma_{c,t,k}}{2}g_{c,t}+\frac{D_{c}\gamma_{c,t,k}}{2}g_{c,k}-\frac{\gamma_{c,t,k}}{2}g_{c,t}g_{c,k}\right).$ (19)
Figure S1: Price and load duration curves for LT and ST models with cross-elasticity terms alongside piecewise-linear own-elastic demand. Panel a) shows the effect of varyign cross-elasticity coefficients $\gamma_{c}$ and $\hat{T}$ on the price duration curve (PDC). Panel b) shows the effect of moving from LT to ST models with cross-elasticity terms on the prices. Panels c) and d) show the load duration curves (LDC) corresponding to the price duration curves shown in panels a) and b), respectively.
[…]
In Fig. S1b, we show that cross-elastic terms can also affect the price identity between LT and ST models with identical capacities. With only own-price elasticity, we have a strict monotonic relationship between price and demand in each hour. As demonstrated in Section 2.3, this leads to a unique solution and identical prices between LT and ST models. When we add cross-price elasticities, demand in each hour depends on prices in multiple hours. This creates, in principle, multiple valid ways to distribute demand across hours while achieving the same total welfare. The LT and ST models can choose different temporal distributions of demand/prices that are equally optimal (i.e. they achieve exactly the same level of utility and supply the same total demand). The results in Fig. S1b for Germany and also Figs. S19 and S20 for Spain and the UK show that, while prices align in the vast majority of hours, the price duration curves are no longer completely identical between LT and ST models.
Appendix C Technology data
@@
H2 turbine 1164.0 €/kWel 5.00 %/a 50.0% 10 a
@@
Appendix D Selection of years in random sequence
The 70 years (1951-2020) we consider are first put in a random sequence using the Fischer-Yates shuffle. The underlying Mersenne Twister random number generator is initialised with the seed 123. For our myopic dispatch analysis in Section 3.3, where an LT model is first optimised on 35 years and then a ST model is run on the remaining 35 years, we take the last 35 years from the shuffled list for the LT model and the first 35 years for the ST model. The resulting years for each model are given in Table S2. For Fig. 3, where the number of years $N$ in the simulation is increased step by step, we use the $N$ last entries in the shuffled list of weather years. That means, for instance, that the years in the 4-year run are included in the 16-year run. The shuffled sequence is not put in chronological order. The resulting years for each case are given in Table S3. All other 70-year runs use the weather years 1951 to 2020 in chronological order. All other single-year runs use 2020 as the weather year.
Model Years (in order)
LT model (35 years) 1960, 1996, 1953, 2020, 1979, 1971, 1998, 2014, 2013, 1989, 1956, 1978, 1951, 2006, 1966, 1995, 2004, 2011, 2009, 1959, 1961, 1954, 2005, 2010, 1972, 1986, 2016, 1975, 1955, 1964, 2019, 2003, 1962, 1985, 1957
ST model (35 years) 2007, 1987, 1974, 1976, 1981, 1993, 1988, 2015, 1958, 2018, 1970, 1990, 1968, 1991, 1965, 1963, 1992, 1973, 2002, 2001, 1982, 1967, 1999, 2017, 1994, 1984, 1977, 1980, 2012, 2000, 1983, 1997, 1969, 1952, 2008
Table S2: Years used in the 35-year LT and ST models.
Number of years Years (in order)
1 year 1957
4 years 2003, 1962, 1985, 1957
16 years 1959, 1961, 1954, 2005, 2010, 1972, 1986, 2016, 1975, 1955, 1964, 2019, 2003, 1962, 1985, 1957
32 years 2020, 1979, 1971, 1998, 2014, 2013, 1989, 1956, 1978, 1951, 2006, 1966, 1995, 2004, 2011, 2009, 1959, 1961, 1954, 2005, 2010, 1972, 1986, 2016, 1975, 1955, 1964, 2019, 2003, 1962, 1985, 1957
48 years 1999, 2017, 1994, 1984, 1977, 1980, 2012, 2000, 1983, 1997, 1969, 1952, 2008, 1960, 1996, 1953, 2020, 1979, 1971, 1998, 2014, 2013, 1989, 1956, 1978, 1951, 2006, 1966, 1995, 2004, 2011, 2009, 1959, 1961, 1954, 2005, 2010, 1972, 1986, 2016, 1975, 1955, 1964, 2019, 2003, 1962, 1985, 1957
70 years 2007, 1987, 1974, 1976, 1981, 1993, 1988, 2015, 1958, 2018, 1970, 1990, 1968, 1991, 1965, 1963, 1992, 1973, 2002, 2001, 1982, 1967, 1999, 2017, 1994, 1984, 1977, 1980, 2012, 2000, 1983, 1997, 1969, 1952, 2008, 1960, 1996, 1953, 2020, 1979, 1971, 1998, 2014, 2013, 1989, 1956, 1978, 1951, 2006, 1966, 1995, 2004, 2011, 2009, 1959, 1961, 1954, 2005, 2010, 1972, 1986, 2016, 1975, 1955, 1964, 2019, 2003, 1962, 1985, 1957
Table S3: Weather years for scenarios with incremental number of years represented.
Appendix E Additional figures